# DreamMix: Decoupling Object Attributes for Enhanced Editability in Customized Image Inpainting

### v1

@@ 5 Conclusion
We present a diffusion-based generative model for subject-driven image inpainting using both text and image guidance. A disentangled inpainting framework is introduced that seamlessly incorporates local subject allocation and global context harmonization to improve object insertion accuracy. Additionally, we propose an attribute decoupling mechanism and a textual attribution substitution module to facilitate text-driven editing of object attributes. Experiments demonstrate the superiority of the proposed method across various subject-driven inpainting applications.
One limitation of our method is that relying solely on text and image guidance may pose challenges to achieving a harmonious interaction with the environment. Incorporating additional conditions (e.g., pose, depth) might be a solution for this issue. Besides, extending our model to support multi-subject inpainting is a future direction.

### v2

@@ 5 Conclusion
We present a diffusion-based generative model for subject-driven image inpainting using both text and image guidance. An attribute decoupling mechanism and a textual attribution substitution module are proposed to facilitate text-driven editing of object attributes. Additionally, a disentangled inpainting framework seamlessly incorporates local subject allocation and global context harmonization to improve object insertion accuracy. Experiments demonstrate the superiority of the proposed method across various subject-driven inpainting applications. One limitation of our method is that relying solely on text and image guidance may pose challenges in achieving a harmonious interaction with the environment. Incorporating additional conditions (e.g., pose, depth) might solve this issue. Besides, extending our model to support multi-subject inpainting is a future direction.
6 Declarations
  • The authors declare no competing interests.
  • Data sharing is not applicable to this article as no new data were created or analyzed in this study.